# Unifying Multitrack Music Arrangement via Reconstruction Fine-Tuning and Efficient Tokenization

### v2

@@ Abstract
Automatic music arrangement streamlines the creation of musical variants for composers and arrangers, reducing reliance on extensive music expertise. However, existing methods suffer from inefficient tokenization, underutilization of pre-trained music language models (LMs), and suboptimal fidelity and coherence in generated arrangements. This paper introduces an efficient multitrack music tokenizer for unconditional and conditional symbolic music generation, along with a unified sequence-to-sequence reconstruction fine-tuning objective for pre-trained music LMs that balances task-specific needs with coherence constraints. Our approach achieves state-of-the-art results on band arrangement, piano reduction, and drum arrangement, surpassing task-specific models in both objective metrics and perceptual quality. Additionally, we demonstrate that generative pretraining significantly contributes to the performance across these arrangement tasks, especially when handling long segments with complex alignment. [^1] Demos available at ae4rtjsyr.github.io. All code will be released upon acceptance.
@@ 1 Introduction
Music arrangement, the art of adapting musical compositions for different performance contexts, plays a fundamental role in music creation by enabling artistic reinterpretation and expanding the accessibility of musical works. It is ubiquitous across various musical domains, from professional music production and education to live performance and digital content creation. However, manual arrangement requires extensive expert knowledge of harmony, instrumentation, and style-specific conventions, making it a highly specialized and demanding process. Automating this task could make arrangement more accessible to individuals with limited musical training and allow composers to focus more on the creative aspects by relieving them of technical groundwork.
Despite the potential benefits, existing methods for automatic arrangement face significant challenges, leading to substantial quality gaps that hinder practical applications. One critical issue lies in tokenization, the foundation of any unconditional or conditional generative music modeling. The commonly adopted REMI tokenizer Huang and Yang (2020) often results in excessively long sequence lengths, a known bottleneck for symbolic music generation Hsiao et al. (2021). Moreover, as an REMI’s extension to multitrack music, REMI+ von Rütte et al. (2023) contains redundant information and suboptimal design choices that lead to the content fragmentation issue in the tokenized sequences, burdening both generative modeling and arrangement performance.
Another limitation of prior arrangement methods is their reliance on task-specific model structures. While generative pre-training has demonstrated great potential in related fields such as NLP, integrating similar approaches into these highly specialized structures is not straightforward. Consequently, these methods fail to fully leverage the generative power of pre-trained music language models (LMs). Given the scarcity of annotated music data, this limitation further constrains their effectiveness.
Finally, existing models often struggle with fidelity and coherence. Specialized arrangement models Zhao et al. (2023, 2024) frequently fail to preserve melody and texture, retaining only harmony, and thus are perceived by users as lacking resemblance to the original piece. Also, our experiments show that even the state-of-the-art piano reduction model Terao et al. (2023) exhibit suboptimal arrangement coherence since different musical segments are handled independently.
To address these limitations, we propose a novel framework for automatic symbolic music arrangement. First, we design a more efficient tokenizer to solve the content fragmentation issue and reduce sequence length and redundancy. Second, we explore the potential of sequence-to-sequence fine-tuning with our pre-trained music LM, enabling task specific models to harness the power of generative pre-training. This approach allows flexible adaptation to various arrangement tasks with limited annotated data. Finally, we validate our unified methodology across multiple tasks, including band arrangement, piano reduction, and drum arrangement.
Our key contributions are summarized as follows:
  • We propose an efficient tokenizer for multitrack music that reduces sequence length and redundant information, improves effectiveness for unconditional pre-training, enhances arrangement performance, and potentially extensible to general music generation tasks.
  • We introduce a self-supervised reconstruction objective to effectively adapt pre-trained music LMs to various music arrangement tasks, ensuring both task-specific requirements and coherence.
  • We achieve state-of-the-art performance on band arrangement, piano reduction, and drum arrangement tasks, surpassing task-specific models in both objective and subjective evaluations.
  • Through ablation studies, we demonstrate how generative pre-training significantly improves bar-to-bar transformations and plays a crucial role in processing longer segments, where models without pre-training fail.
@@ 2.1 Automatic Arrangement in Multitrack Symbolic Music
Music arrangement encompasses various tasks like chord progression creation Yi et al. (2022), lead sheet orchestration Wang et al. (2024), and composition adaptation Zhao et al. (2024). This paper focuses on adapting existing music to new instruments or ensembles, which is crucial for multi-style music generation and style transfer.
Earlier supervised approaches trained end-to-end models using parallel data (e.g., piano-to-orchestra Crestel and Esling (2016) or band-to-piano Terao et al. (2022, 2023)), which are extremely challenging to prepare and inherently restrict the direction of arrangement. Classification-based methods such as Dong et al. (2021) attempt to predict instrument labels for each note, but they lack creativity by generating no new content and suffer from inflexibility, being incapable of handling unseen instrument combinations. Recent self-supervised methods like Q&A Zhao et al. (2023, 2024) offer more flexible solutions without requiring parallel datasets or fixed instrument set assumptions. However, Zhao et al. (2023) assumes prior knowledge of track-wise output distributions (i.e., note density and pitch histograms), which is impractical for arrangement tasks since such knowledge is unavailable before generation. While Zhao et al. (2024) attempts to solve this by introducing a separate prior model to estimate the track-wise distributions, this approach comes at the cost of compromised arrangement fidelity. Additionally, infilling-based models such as Composer’s Assistant Malandro (2023, 2024), although capable of handling additive arrangement scenarios such as drum arrangement, do not preserve the contents of the original composition in the generation output and are thus fundamentally unsuited for band arrangement, where the primary aim is to reinterpret existing musical content.
2.2 Generative Symbolic Music Models
Transformer networks have advanced music generation tasks like chord-to-melody generation Madaghiele et al. (2021), accompaniment arrangement Ren et al. (2020), and music inpainting Malandro (2023, 2024). However, these models remain task-specific, with limited exploration of unconditional pre-training and task-specific fine-tuning. Regarding controlled generation, existing works primarily focus on using simple high-level attributes (style Choi et al. (2020); Lu et al. (2023), structure Zhang et al. (2022), and sentiment Makris et al. (2021)), with minimal investigation of music-based content conditioning. As large symbolic music LMs emerge Dong et al. (2023); Qu et al. (2024); Wu et al. (2024), it becomes crucial to explore more complex control signals, such as music itself, and expand their application to broader scenarios.
2.3 Symbolic Music Tokenizers for Transformers
@@
Transformer-based music modeling requires converting data into token sequences, which is challenging due to music’s multi-stream nature. Many tokenizers treat music as a single stream using MIDI-like note-event encoding, representing each musical note with multiple tokens (onset time, pitch, duration, velocity, instrument). Some use absolute timing Huang et al. (2018); Gardner et al. (2021); Zeng et al. (2021); Ens and Pasquier (2020), while others use metrical durations Ren et al. (2020); Huang and Yang (2020); von Rütte et al. (2023). Among them, the REMI Huang and Yang (2020) tokenizer is widely adopted for single-track music and has been extended to multitrack in von Rütte et al. (2023) by adding an instrument token as an attribute of each note.
However, such extension, along with all approaches that strictly order notes temporally for multitrack music, has an inherent limitation—content fragmentation, i.e., notes from one instrument being interleaved with notes from other instruments playing simultaneously. As shown in Figure 1(a), the content of instrument i-29 (distorted electric guitar, highlighted in orange) spreads across the entire sequence, interleaved by notes from instrument i-80 (synth lead, highlighted in red). Consequently, tokenization results can be highly inconsistent for same instrument-wise contents once its context (interleaving contents) differs, bringing extra difficulty for generative modeling.
2.4 Prompt-based Fine-Tuning for Conditional Generation
Prompt-based fine-tuning in NLP Liu et al. (2023) uses control tokens to guide pre-trained LM outputs, enabling attribute control over style Sennrich et al. (2015), length Kikuchi et al. (2016), and pronunciation Ou et al. (2023). Combining such strategy with LMs pre-trained with standard left-to-right language modeling objective Radford et al. (2019); Brown et al. (2020) can effectively generates fluent, condition-aligned outputs, providing insights for the music arrangement problems, where similar pre-training and control mechanisms can potentially generate music that is musically appealing and meets task-specific requirements.
@@ 3 Method
To bridge the gaps in previous research, we adopt a pre-train and fine-tune paradigm for the arrangement task, design an efficient multitrack music representation for Transformer networks that makes both pre-training and fine-tuning more effective, and develop a unified objective for music-to-music transformation that addresses both task-specific requirements and coherence. We test the effectiveness of this approach on multiple typical music arrangement tasks.
@@
Figure 1: REMI-z and REMI+ tokenization for a same bar. Content of the same instruments are highlighted with same color.
(a) Detailed settings for each arrangement task. The symbol $\bigoplus$ denotes concatenation of component sequences.
(b) An example of the tokenized sequence for a band arrangement task. Special tokens $<$$\tt{SEP}$$>$, $<$$\tt{INSTRUMENT}$$>$, $<$$\tt{CONTENT}$$>$, and $<$$\tt{HISTORY}$$>$ are used to separate different components. Notes of each instrument are highlighted with different colors.
Figure 2: Music segments are decomposed into three subsequences: instruments, content, and target-side history. These components form the condition sequence, with the original music as the target sequence. The model is trained to reconstruct the music from these components.
3.1 REMI-z Representation
Meaning Token X’s range
Instrument type i-X 0$\sim$128
Note’s within-bar position o-X 0$\sim$127
Note’s pitch (non-drum) p-X 0$\sim$127
Note’s pitch (drum) p-X 128$\sim$255
Note’s duration d-X 0$\sim$127
End of a bar b-1 -
Time signature s-X 0$\sim$253
Tempo t-X 0$\sim$48
Table 1: REMI-z vocabulary with pitch token distinctions
Instruments have inherent syntactic rules governed by their physical properties and conventional playing patterns. A good arrangement, or in broader terms, any musical composition, must ensure each instrumental part is both executable and idiomatic. However, as mentioned earlier, time-ordered tokenizers struggle with instrument-wise content fragmentation, making it difficult to model these syntactic rules. To address this issue, we propose REMI-z, a tokenization scheme that does not sort notes strictly by time but instead prioritizes the sequence continuity of content for each instrument.
The vocabulary of REMI-z is shown in Table 1, similar to that of REMI+ von Rütte et al. (2023) (detailed in Appendix A). The key innovation lies in how notes are organized. REMI-z tokenize arbitrary-length MIDI files into a list of bar sequences, each ending with a special end-of-bar token (b-1). Within each bar, notes are further grouped into track sequences, where each track corresponds to a specific instrument. Each track sequence begins with an instrument token specifying the instrument type, followed by all notes played by that instrument within the bar. Track sequences are ordered by their average pitch (from high to low), and within each track sequence, notes are represented as triplets of (position, pitch, duration). These notes are sorted first by position (left to right) and then by pitch (high to low) before being flattened into a 1-D sequence. If two adjacent notes share the same position, the second position token is omitted.
The primary difference between REMI-z and REMI+ is that in REMI-z, notes belonging to the same instrument are grouped together within each bar (as in Figure 1(b)), significantly alleviating the content fragmentation issue and facilitating instrument-specific syntactic modeling. Another advantage is that the sequence length is reduced, as instrument tokens appear only once per track sequence in REMI-z, whereas REMI+ needs an instrument token for each note.
We adopt REMI-z tokenization to pre-train a Transformer decoder using a standard next-token-prediction objective on a large unlabeled corpus and then proceed to convert the pre-trained model to task-specific models with the objective in the following section.
3.2 Arrangement Fine-Tuning
Music arrangement involves adapting given musical content to a new set of instrumental constraints. With a well-trained content- and instrument-controllable generation model, novel arrangements can be generated from the original composition by assigning new instrumentation constraints at inference time. To train this capability without relying on parallel data, such as multiple versions of the same song arranged for different instrumentations, we propose a context-aware self-supervised segment-level reconstruction objective. Let $y^{(t)}$ denote the $t$-th segment of a music piece. The fine-tuning objective is defined as:
$\mathcal{L}(\theta)=-\log\,p_{\theta}\big{(}$ $\mathcal{T}_{\text{task}}(y^{(t)})\,\big{|}\,\mathrm{I}(\mathcal{T}_{\text{task}}(y^{(t)})),$
$\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)})),\,\mathcal{T}_{\text{task}}(y^{(t-1)})\big{)},$ (1)
where $\theta$ represents the model parameters, $\mathrm{I}(\cdot)$ and $\mathrm{C}(\cdot)$ extract instrument and content information, $\mathcal{S}_{\text{task}}$ and $\mathcal{T}_{\text{task}}$ represent selecting task-specific source and target tracks, and $y^{(t-1)}$ provides target-side history. Equation 3.2 is implemented with a standard LM objective to the sequence [condition]<sep>[target], as illustrated in Figure 2(b), with cross-entropy loss computed only on the target subsequence.
During training, the instrument, content, and history conditions are derived from the same music being reconstructed. The model is trained to reconstruct the original music, learning to interpret the given musical content with a specific instrument set in a certain context. During inference, users can flexibly define instrument constraints, while the content comes from the music to be arranged, and history from the model’s previously generated outputs.
The instrument condition, derived from $\mathrm{I}(\cdot)$, specifies the desired instruments to be used in the segment. During training, it includes all instrument tokens present in the target-side sequence. Additionally, the relative positions of tokens within this sequence encode important information about the voice relationships between instruments: tokens earlier in the sequence correspond to instruments with higher average pitches, and vice versa. At inference, users have the flexibility to specify both the instruments and their voice relationships for each segment.
The content condition from $\mathrm{C}(\cdot)$ represents the musical notes from the original composition, encoded as a content sequence. This sequence provides precise information about what notes are played and their timing, but without specifying instruments. To construct the content sequence, the following steps are applied: (1) remove all instrument tokens from REMI-z, (2) sort notes strictly by their temporal positions, (3) sort notes within the same position by pitch in descending order, and (4) merge duplicate notes with identical positions and pitches. This representation ensures a clean, time-ordered view of the musical content of a multitrack segment without any instrument-related information.
Lastly, we introduce a history condition $\mathcal{T}_{\text{task}}(y^{(t-1)})$ that encourages contextual coherence between segments during the arrangement process to tackle the style inconsistency issue brought by segment-level processing, inspired by the success of historical context in machine translation Tiedemann and Scherrer (2017). During training, the history is provided through teacher-forcing using the ground truth REMI-z sequence from the previous segment’s target sequence, while at inference time, it is generated autoregressively based on the model’s output from the preceding segment. This mechanism ensures that each segment is generated with awareness of its musical context, leading to coherent arrangements.
3.3 Tasks
We evaluate our method on the following tasks, which are representative of typical arrangement scenarios in music composition studies. Each task assesses different aspects of model capabilities.
3.3.1 Band Arrangement
This task involves arranging an existing piece of music for arbitrary combinations of instruments. The model must understand the properties and typical playing styles of each instrument to allocate or generate new notes appropriately. In this setup, both $\operatorname{\mathcal{S}_{\text{task}}}$ and $\operatorname{\mathcal{T}_{\text{task}}}$ are defined as identity transformations, directly returning the original REMI-z sequence, i.e., $\operatorname{\mathcal{S}_{\text{task}}}(y)=\operatorname{\mathcal{T}_{\text{task}}}(y)=y$. Note that we focuses exclusively on pitched instruments and pre-remove drum tracks from $y$ due to REMI-z’s drum-specific pitch tokens. Drum arrangement is addressed as a separate task.
To foster creativity, we randomly delete a small proportion of tracks in $\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)}))$ during training, encouraging the model to generate new notes that are compatible with the musical context but absent from the input music during inference. Special care is taken to preserve the melody during random deletions by ensuring that the track with the highest average pitch remains untouched. Additionally, duration tokens are removed from the content sequence to allow the model to generate durations that best suit the playing styles of the desired instruments, rather than merely copying from inputs. The segment length is set to 1 bar.
3.3.2 Piano Reduction

### v3

# Unifying Symbolic Music Arrangement: Track-Aware Reconstruction and Structured Tokenization

@@ Abstract
We present a unified framework for automatic multitrack music arrangement that enables a single pre-trained symbolic music model to handle diverse arrangement scenarios, including reinterpretation, simplification, and additive generation. At its core is a segment-level reconstruction objective operating on token-level disentangled content and style, allowing for flexible any-to-any instrumentation transformations at inference time. To support track-wise modeling, we introduce REMI-z, a structured tokenization scheme for multitrack symbolic music that enhances modeling efficiency and effectiveness for both arrangement tasks and unconditional generation. Our method outperforms task-specific state-of-the-art models on representative tasks in different arrangement scenarios—band arrangement, piano reduction, and drum arrangement, in both objective metrics and perceptual evaluations. Taken together, our framework demonstrates strong generality and suggests broader applicability in symbolic music-to-music transformation. [^1] Demos available at ae4rtjsyr.github.io. All code will be released upon acceptance.
@@ 1 Introduction
Music arrangement is the art of adapting compositions for performance contexts that differ from their original forms corozine2015arranging . It plays a central role in many music creation process, including professional production, live performance, music education, and digital content creation. Automating this process can expand music accessibility and accelerate music creation. Although arrangement forms vary—e.g., rewriting for different instruments (reinterpretation) dong2021towards , simplifying for solo performance (reduction) terao2022difficulty ; terao2023neural , or adding new tracks (additive generation) malandro2023composer ; ComposersAssistant2 —they share a common structure: generating new music tracks conditioned on existing ones under explicit content and instrument constraints. However, prior work typically addresses each task independently zhao2023accomontage ; terao2023neural ; terao2022difficulty ; ComposersAssistant2 , using specialized model architectures and training schemes. Such designs lack cross-task generality, increase implementation cost, and fail to leverage the musical knowledge learned by large-scale pre-trained symbolic models that could potentially improve the arrangement quality.
Meanwhile, generative modeling of symbolic music, i.e., music in notation-based formats, has advanced rapidly with large-scale pre-trained models that capture rich musical styles and structures via autoregressive modeling lu2023musecoco ; dong2023multitrack ; qu2024mupt ; wu2024melodyt5 ; wang2025notagen ; pasquier2025midi . Inspired by natural language processing, these models scale to billions of parameters and are trained on vast corpora. Although their unconditional generation quality has improved substantially, their applications in real-world conditional generation remains relatively limited. Prior work primarily targets coarse-level control, such as style choi2020encoding ; lu2023musecoco , structure zhang2022structure , polyphony level pasquier2025midi , or sentiment makris2021generating , while fine-grained conditioning on existing musical content—precisely what arrangement tasks demand—remains underexplored.
Figure 1: Overview of the proposed unified framework for each arrangement task. The symbol $\bigoplus$ denotes concatenation of component sequences. Music segments are decomposed into three subsequences: instruments, content, and target-side history. These components form the condition sequence, with the relevant tracks from the original music as the target sequence. The model is trained to reconstruct the music from these components.
This underexploration is likely not due to the absence of pre-trained models or their representational power, but to the lack of training objectives that support flexible adaptation to the diverse requirements of arrangement tasks. While sequence-to-sequence fine-tuning is conceptually a natural solution, it relies on parallel datasets—collections of the same piece arranged for different instrumentations (e.g., orchestra and piano)—which are extremely scarce. Even when adopting some level of internal parallel data terao2022difficulty ; terao2023neural , these datasets constrain output directions to fixed mappings (e.g., orchestra to piano). In contrast, real-world arrangement demands far greater flexibility: the ability to transform arbitrary input instrumentations into arbitrary targets—i.e., support any-to-any transformations.
To address these limitations, we propose a unified framework for symbolic music arrangement that enables a pre-trained symbolic generative model to be fine-tuned across diverse tasks through a single self-supervised training pipeline (Figure 1). This facilitates the transfer of musical knowledge from generative pre-training to improve arrangement quality, while reducing the need for task-specific model design. The core is a context-aware, segment-level reconstruction objective: the model reconstructs multitrack music from its disentangled components: content (notes executed) and style (instrumentation). We further observe that strictly time-ordered tokenizations (e.g., REMI+ von2023figaro ) introduce redundancy and fragment track content, hindering instrument-level control and modeling. To address this, we propose a structured tokenization scheme that relaxes global time ordering while promoting track-wise continuity, enabling more consistent encoding across musical contexts and better learning under the proposed objective. Our main contributions are as follows:
  • We propose a unified framework for symbolic music arrangement that supports flexible instrumentation transformation across multiple typical arrangement scenarios, all without requiring parallel data. Central to our approach is a shared reconstruction objective defined over token-level disentangled note properties, which enables a single generative model to adapt through lightweight fine-tuning.
  • We design an efficient and modeling-friendly tokenization scheme for multitrack music that produces shorter sequences with lower complexity, facilitates instrument-level control and modeling that are crucial for arrangement performance. It further reduces note-level perplexity in unconditional generation, indicating potential utility beyond arrangement tasks.
  • Instantiated with a small model and modest-scale pre-training, our system outperforms task-specific SOTA baselines on three representative arrangement tasks, each corresponding to a distinct scenario—band arrangement (reinterpretation), piano reduction (simplification), and drum arrangement (additive generation)—in both objective and subjective evaluations.
@@ 2.1 Automatic Arrangement in Multitrack Symbolic Music
Symbolic music arrangement encompasses a variety of tasks such as chord progression generation yi2022accomontage2 , orchestration from lead sheets wang2024whole , and instrumentation transfer zhao2023accomontage . In this work, we focus on adapting multitrack music to new instrumentations—a representative setting that requires fine-grained control over musical content and instrumentation.
Early supervised learning approaches rely on parallel datasets (e.g., piano-to-orchestra crestel2016live , band-to-piano terao2022difficulty ; terao2023neural ), which are expensive to construct and inherently constrain the direction of arrangement. Classification-based methods dong2021towards approach the arrangement problem by learning to assign instrument labels to each note from a dataset with fixed instrumentation. However, such models are limited in expressiveness, as they cannot modify musical content—e.g., adding, removing, or altering notes—to suit a target instrument combination. Moreover, their reliance on fixed instrument sets hinders generalization to unseen combinations.
Recent self-supervised methods zhao2023q ; zhao2023accomontage avoid these constraints by guiding generation with pre-defined or autoregressively modeled high-level descriptors (e.g., pitch histogram, note density). While effective for maintaining stylistic coherence, they separate content modeling from style realization. As a result, instrument playing styles are either fixed or modeled independently of input variations. Since musical material often changes between sections, this decoupling can degrade fidelity—causing arrangements to fail relect core aspects such as melody or texture, resulting in noticeable perceptual divergence. This suggests a more fidelity-oriented approach: integrating style modeling into the generation process to allow dynamic adaptation to input content.
Additionally, infilling-based models such as Composer’s Assistant malandro2023composer ; ComposersAssistant2 , although capable of handling additive generation scenarios such as drum arrangement, do not preserve the music essence of the original composition and are thus unsuited for reinterpretation and simplification.
2.2 Symbolic Music Tokenization
@@
Transformer-based symbolic music modeling requires converting musical data into token sequences, a nontrivial task due to music’s inherent multi-stream structure, i.e., multiple instruments playing concurrently. Many existing tokenization schemes adopt a linearized, note-event-based encoding where each musical note is represented as a tuple of attribute tokens (e.g., onset, pitch, duration, velocity, instrument). Some use absolute timing Huang2018MusicTG ; Gardner2021MT3MM ; Zeng2021MusicBERTSM ; Ens2020MMME , while others use metric durations ren2020popmag ; huang2020pop ; von2023figaro . Among them, the REMI representation huang2020pop , originally designed for single-track music, has been extended to REMI+ von2023figaro for flexible tokenization of multitrack music, by associating each note with an instrument token.
However, existing time-ordered tokenization schemes such as REMI+ face structural limitations that hinder track-wise control and modeling. By flattening multitrack music into a strictly time-ordered sequence, REMI+ interleaves events from different instruments, resulting in content fragmentation. As illustrated in Figure 3(a), notes from the same instrument (e.g., i-29, distorted electric guitar, in orange) are interrupted by those from concurrent instruments (e.g., i-80, synth lead, in red). This leads to two major issues. First, the lack of structured syntax makes it difficult to delineate track boundaries, hindering enforcement of user-specified instrument constraints and can result in spurious instruments in arrangement ouptuts. Second, identical track-wise content may be tokenized differently depending on concurrent instrument activity. This context sensitivity reduces the repetition of typical per-instrument patterns in the training data—a key factor for learning accurate “instrument syntax” determined by physical constraints and idiomatic playing patterns of each instrument, important for arrangement quality. While some recent work explores vocabulary-level compression (e.g., Byte Pair Encoding fradet-etal-2023-byte ) to reduce sequence length—an approach orthogonal and complementary to tokenization scheme design—it does not address the structural issues discussed above.
@@ 3 Method
Our goal is to fine-tune a single pre-trained symbolic music generative model across diverse arrangement tasks via a unified pipeline. The core is a segment-level reconstruction objective (§3.1) over token-level disentangled representations of style (instrumentation) and content (musical notes). To support track-wise control and modeling, we introduce a structured tokenization scheme (§3.2) that reduces cross-instrument fragmentation and enhances learning under this objective.
3.1 Reconstruction from Token-Level Disentangled Multi-Streams with Context Awareness
Symbolic music offers a unique opportunity for token-level disentanglement: each note is represented as a list of semantically independent tokens, each describing a unique property of the note (e.g., onset, duration, pitch, instrument), allowing content (what is played) and instrumentation (by whom it is played) to be explicitly separated. This level of structural redundancy is rare in natural language, where sub-word tokens are atomic.
Building on this observation, we formulate arrangement as a self-supervised, segment-level reconstruction task. The input music is decomposed into three token streams—instrument, content, and preceding context—and a pre-trained symbolic music generative model is fine-tuned to reconstruct the desired tracks of the multitrack music from these components.
Instrument condition. The instrument condition specifies the desired instruments to be used in the segment. During training, it includes all instrument tokens from the target sequence. Their order encodes pitch register relationships across tracks: instruments with higher average pitch are placed earlier. At inference, users can specify arbitrary desired instruments (instrument control) and define their relative pitch register ordering (voice control).
Content condition. The content condition is derived from the original composition, encoded as a content sequence without instrument information. To obtain it, we remove all instrument tokens from the original multitrack sequence, sort notes by onset time, then by descending pitch, and merge duplicates. This yields a clean time-ordered note sequence conveying what is played, independent of how or by whom.
History condition. The history condition provides musical context from the preceding segment, encouraging inter-segment coherence at inference time. During training, it is provided via teacher-forcing with the complete tokenized previous segment; at inference, it is autoregressively generated. This mechanism helps maintain coherent arrangement style across segments, especially important for long-form (e.g., song-level) arrangement scenarios.
Let $y^{(t)}$ denote the $t$-th segment of a music piece. The fine-tuning objective is defined as:
$\mathcal{L}(\theta)=-\log\,p_{\theta}\big($ $\mathcal{T}_{\text{task}}(y^{(t)})\,\big|\,\mathrm{I}(\mathcal{T}_{\text{task}}(y^{(t)})),\,\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)})),\,\mathcal{T}_{\text{task}}(y^{(t-1)})\big),$ (1)
where $\theta$ represents the model parameters, $\mathrm{I}(\cdot)$ and $\mathrm{C}(\cdot)$ extract instrument and content conditions, $\mathcal{S}_{\text{task}}$ and $\mathcal{T}_{\text{task}}$ are filters that select task-specific source and target tracks respectively, and $y^{(t-1)}$ provides target-side history. Equation 1 is implemented using a standard next-token prediction objective on the sequence [condition]<sep>[target] (Figure 2), with cross-entropy loss computed only on the target subsequence .
During training, the instrument, content, and history conditions are derived from the same music being reconstructed. The model thus learns to reinterpret musical content with various instrument combinations under specific contexts, enabling diverse arrangement behaviors.
Figure 2: An example of the tokenized sequence for the band arrangement task. Special tokens $\tt{SEP}$>>, $\tt{INSTRUMENT}$>>, $\tt{CONTENT}$>>, and $\tt{HISTORY}$>> are used to separate different components. Tokens starting with o-, i-, p-, d- represents the onset, instrument ID, pitch, and duration of notes.
3.2 REMI-z: Tokenizing Multitrack Music with Track-Wise Continuity
@@
Figure 3: REMI+ and REMI-z tokenization for the same bar. Contents of the same instruments are highlighted with the same color. See Appendix A for complete vocabulary.
As discussed in Section 2.2, strictly time-ordered tokenization schemes suffer from content fragmentation, a structural limitation that hinders arrangement performance. To mitigate this issue, we propose REMI-z, a tokenization scheme that heuristically prioritizes track-wise continuity over global time ordering. Specifically, REMI-z processes MIDI files into a list of bar sequences, each containing multiple track sequences, where each track corresponds to a unique instrument. Within each track, note events are sorted by their onset position, then by descending pitch, and grouped under a single instrument token. Tracks within a bar are then ordered by their average pitch (high to low), forming a bar-level token sequence that ends with a special end-of-bar token. The complete vocabulary and tokenization examples are shown in Appendix A.
This zig-zag organization—reflected in the REMI-z name—offers several modeling advantages: (1) instrument-wise content is locally contiguous (Figure 3(b)), reducing content fragmentation and enabling clear track boundaries; (2) sequence length is reduced by eliminating redundant instrument tokens; (3) temporal structure is preserved both within individual track sequences and between bars.
We adopt REMI-z to tokenize all data for both pre-training and fine-tuning. Beyond improving arrangement performance, we also observe that REMI-z produces sequences with lower information entropy and, when used for unconditional generative training, leads to better note-level modeling (see § 5.4). This suggests that its track-continuity design not only benefits arrangement tasks, but may also support general symbolic music modeling.
3.3 Task Instantiations
We evaluate our method on three representative music arrangement tasks that reflect typical arrangement scenarios, each assessing different capabilities of the model: band arrangement (reinterpretation), piano reduction (simplification), and drum arrangement (additive generation).
Band Arrangement. This task evaluates the model’s ability to reinterpret an existing piece using arbitrary combinations of pitched instruments. The model must learn the properties and idiomatic playing styles of various instruments to reallocate or generate notes appropriately. Both $\operatorname{\mathcal{S}_{\text{task}}}$ and $\operatorname{\mathcal{T}_{\text{task}}}$ are identity mappings (i.e., $\operatorname{\mathcal{S}_{\text{task}}}(y)=\operatorname{\mathcal{T}_{\text{task}}}(y)=y$), and drum tracks are removed from the input. To encourage creative rewriting, we randomly remove a subset of tracks from the content condition $\mathrm{C}(\mathcal{S}_{\text{task}}(y^{(t)}))$ during training, while ensuring the melody’s content is preserved (detailed in Appendix B.2). Duration tokens are also removed from the content stream, allowing the model to infer track-specific note durations, i.e., articulations, suitable for interpreting the content under desired instrumentations. Segment length is set to 1 bar.